OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (jdoe0001):

what's the slope of that line?

OpenStudy (anonymous):

2-1 over 5-4 =1 over 1 which equals 1 y-1 = 1(x-4) y-1=1x-4 y=-3